[Hallc_running] [Revised Logentry] Hall C RC daily update

simona at jlab.org simona at jlab.org
Mon Aug 3 16:00:06 EDT 2020

Logentry Text:
Yesterday we decided to go ahead and give it a try at imaging the Carbon hole although the beam looked kind of fat when looking at the H07 harps. Beam was centered on C-hole and the reference cell during the owl shift. Ion chamber calibrations were done for "no target" and the C target. The MCC crew stopped short of performing the IC calibration on the reference cell for reasons unknown. 

This morning while waiting for the Moller experts to start a Moller measurement we requested a BCM calibration. We didn't get too far with that: MCC could not deliver stable beam at 30 muA (they tried for about an hour). We aborted the BCM calibration run as it was time to do a Moller measurement. Then storms hit Newport News which lead to the usual RF instabilities. We are still waiting for beam to resume the Moller measurement.
After the Moller we will do an IC calibration on the reference cell (I got assurance from the day crew chief that it won't be a problem this time) and then we will probably move into production on glass cells.

We will have a RC meeting today at 4:15 pm. This is how to connect via bluejeans: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3820141
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200803/70e8ba7d/attachment.html>

More information about the Hallc_running mailing list